Developing Interactive PVM-based Parallel Programs on Distributed

Computing Systems within AVS Framework

Gang Cheng, Geoffrey C. Fox, Kim Mills and Marek Podgorny

Northeast Parallel Architectures Center Syracuse University, Syracuse, NY 13244-4100

Contact Person: Gang Cheng Tel: 315-443-2083(o), 315-479-8767(h) Fax: 315-443-1973 Email: gcheng@npac.syr.edu

Paper Track: Developing Applications with AVS Inc. Products Presenter: Gang Cheng